OpenStudy (anonymous):

How many perfect cubes divide evenly into 13,824? Please explain.

OpenStudy (asnaseer):

First step would be to express 13824 and the product of primes - can you do that?

OpenStudy (anonymous):

So, you're supposed to multiply prime numbers to get 13824?

OpenStudy (asnaseer):

yes, e.g.:\[18=2\times3^2\]

OpenStudy (asnaseer):

any number can be expressed as the product of powers of prime numbers

OpenStudy (asnaseer):

its called the prime factorisation of a number

OpenStudy (anonymous):

What are you supposed to do after that?

OpenStudy (asnaseer):

lets take a smaller example - 216. This can be written as:\[216=2^3\times3^3\]Can you see now how to get all perfect cubes that divide evenly into 216?

OpenStudy (anonymous):

Yes.

OpenStudy (asnaseer):

So try the same thing with the number you are given

OpenStudy (anonymous):

Okay, but one of the things confusing me is that the answer is supposed to be 7.

OpenStudy (asnaseer):

what prime factorisation did you get for 13824?

OpenStudy (anonymous):

I got 2\[2^{9} and 3^{3}\]

OpenStudy (asnaseer):

correct, so:\[13824=2^9\times3^3\]Now you need to work out how many cubes can be formed from this - what do you get?

OpenStudy (anonymous):

What do you mean by how many cubes?

OpenStudy (asnaseer):

e.g. \(3^3\) is a perfect cube that would divide evenly into 13824

OpenStudy (anonymous):

Oh, so for 2^9, would you break that into 2 further cubes that would be 2^3 and 2^3?

OpenStudy (asnaseer):

well from \(2^9\) you can see that you have these perfect cubes: \(2^3\), \((2^2)^3=4^3\)

OpenStudy (asnaseer):

similarly, you can use combinations of 2 and 3 to get other cubes

OpenStudy (asnaseer):

and of course \(2^9\) itself is also a perfect cube

OpenStudy (anonymous):

So, for this question then, you have 2^3, 2^3 and 3^3, do you just add the numbers then like, 2,2, and 3?

OpenStudy (asnaseer):

no

OpenStudy (asnaseer):

\(2^3\) is one cube that divides evenly into 13824 \((2^2)^3=4^3\) is another cube that divides evenly into 13824 \((2^3)^3=8^3\) is another cube that divides evenly into 13824 \(3^3\) is another cube that divides evenly into 13824 now find the others that can be formed by combining the two prime factors 2 and 3 to get perfect cubes

OpenStudy (anonymous):

I'm not sure...

OpenStudy (asnaseer):

HINT: One of them would be \((2\times3)^3=6^3\)

OpenStudy (anonymous):

I'm sorry, but I still don't understand how to to solve it. I get what you're saying with prime factorization and stuff, but I don't get how to find the answer.

OpenStudy (asnaseer):

np - I must be explaining it badly - let me try another approach...

OpenStudy (asnaseer):

we know that:\[13824=2^9\times3^3\]and we are asked to find all perfect cubes that divide evenly into this number. So, we are supposed to find cubes, \(x^3\), such that:\[13824=x^3\times\text{some other number}\]does that make sense so far?

OpenStudy (anonymous):

Yupp.

OpenStudy (asnaseer):

good - so a simple one would be:\[13824=2^3\times(2^6\times3^3)\]giving us \(x^3=2^3\) make sense?

OpenStudy (anonymous):

Right.

OpenStudy (asnaseer):

so basically we are being asked to work out in how many different ways can we write the product \(2^9\times3^3\) such that it is the product of a cube and some other number

OpenStudy (asnaseer):

the examples we have found so far are: